# Stochastic Adaptive Estimation in Polynomial Curvature Shape State Space for Continuum Robots

### v4

@@ VI-C Evaluation Metrics
In the Cartesian space, we consider two common error metrics: tip error and shape error[9]. At any point along the continuum robot’s arc, we define two error metrics. The position error $e_{p}$ is expressed as a percentage relative to the arc length up to that point, while the rotation error $e_{r}$ is given as the relative error in the bending angle per 100 $mm$ of arc length. Specifically, we define:
@@
The tip errors are given directly by (70) and (71) using corresponding data at the tip. The shape errors are the average errors across multiple evenly distributed locations along the arc length, including the tip, similar to the metric used in [60]. Here we compute shape errors using 10 sampled locations along the arclength.
@@ VII-B Data Collection and Processing
The experimental validation employed four distinct scenarios, comprising three single-segment bending cases (Fig. 14) and one two-segment bending case (left panel of Fig. 16). Scenario I: free bending from an initially straight configuration (no external load). Scenario II: tip unloading—the tip initially carries a 50 g mass that is placed onto a fixed platform during the motion. Scenario III: distributed loading—a wire threaded through the last ten disks is pulled downward by a 100 g mass. Scenario IV: two-segment bending with independent tip loads—100 g at the first-segment tip and 50 g at the second-segment tip. All force loading(unloading) was performed by manually adding(removing) weights in a quasi-static manner. Each loading(unloading) cycle included a brief ramp-up(down) (around 2 s), followed by a constant-load phase of a few seconds corresponding to the continuum robot motion time. Following our simulation methods, changes in the bending angle were converted into motor commands to actuate the wire consistently using inverse kinematics of joint-to-configuration space, as detailed in the supplementary Appendix A. During each scenario, we collected raw pose data from all Aruco markers throughout the bending process within a ROS2 environment[61], enabling us to compute disk poses relative to the base frame. Since our study focuses solely on in-plane bending, determining the bending plane was essential. We collected all y-axes from the Frames {Gs} on the disks and employed Singular Value Decomposition (SVD) to pinpoint the principal y-axis. This axis helped establish the normal to the bending plane. Using this principal y-axis alongside the origin of the base frame, we defined the bending plane, enabling the projection of the 6 DOF pose data into 3 DOF within this plane. The projected poses from the Aruco marker at the tip detected by camera_m served as the measurement data for our analyses.
@@ VII-D Results and Discussion
Finally, with respect to existing literature, direct quantitative comparisons are hindered by differing experimental conditions (e.g., sensor noise, hardware design, or force profiles) that heavily affect estimation outcomes. Nonetheless, the Constant-Curvature EKF ($M^{\mathit{(0)}}$) remains a widely accepted baseline[8, 11, 13], and 1st-order polynomial curvature ($M^{\mathit{(1)}}$) aligns with popular linear-curvature assumptions (e.g., Euler curves[40] or Clothoid curves[62]). Hence, assessing our IMM-EKFs under identical conditions against these standard single-model EKFs demonstrates the measurable advantages and robustness offered by the proposed stochastic observer in polynomial curvature space. The ability to flexibly employ higher-order models ($M^{\mathit{(2)}}$) further underscores its potential for improved accuracy when the application demands richer shape representations.
@@ VIII Conclusion
An extended Supplementary Appendix (Appendices A–C) is available as an ancillary file with the arXiv version of this article[63].

### v5

@@ VI-C Evaluation Metrics
In the Cartesian space, we consider two common error metrics: tip error and shape error[shi2016shape]. At any point along the continuum robot’s arc, we define two error metrics. The position error $e_{p}$ is expressed as a percentage relative to the arc length up to that point, while the rotation error $e_{r}$ is given as the relative error in the bending angle per 100 $mm$ of arc length. Specifically, we define:
@@
The tip errors are given directly by (70) and (71) using corresponding data at the tip. The shape errors are the average errors across multiple evenly distributed locations along the arc length, including the tip, similar to the metric used in [rao2021using]. Here we compute shape errors using 10 sampled locations along the arclength.
@@ VII-B Data Collection and Processing
The experimental validation employed four distinct scenarios, comprising three single-segment bending cases (Fig. 14) and one two-segment bending case (left panel of Fig. 16). Scenario I: free bending from an initially straight configuration (no external load). Scenario II: tip unloading—the tip initially carries a 50 g mass that is placed onto a fixed platform during the motion. Scenario III: distributed loading—a wire threaded through the last ten disks is pulled downward by a 100 g mass. Scenario IV: two-segment bending with independent tip loads—100 g at the first-segment tip and 50 g at the second-segment tip. All force loading(unloading) was performed by manually adding(removing) weights in a quasi-static manner. Each loading(unloading) cycle included a brief ramp-up(down) (around 2 s), followed by a constant-load phase of a few seconds corresponding to the continuum robot motion time. Following our simulation methods, changes in the bending angle were converted into motor commands to actuate the wire consistently using inverse kinematics of joint-to-configuration space, as detailed in the supplementary Appendix A. During each scenario, we collected raw pose data from all Aruco markers throughout the bending process within a ROS2 environment[doi:10.1126/scirobotics.abm6074], enabling us to compute disk poses relative to the base frame. Since our study focuses solely on in-plane bending, determining the bending plane was essential. We collected all y-axes from the Frames {Gs} on the disks and employed Singular Value Decomposition (SVD) to pinpoint the principal y-axis. This axis helped establish the normal to the bending plane. Using this principal y-axis alongside the origin of the base frame, we defined the bending plane, enabling the projection of the 6 DOF pose data into 3 DOF within this plane. The projected poses from the Aruco marker at the tip detected by camera_m served as the measurement data for our analyses.
@@ VII-D Results and Discussion
Finally, with respect to existing literature, direct quantitative comparisons are hindered by differing experimental conditions (e.g., sensor noise, hardware design, or force profiles) that heavily affect estimation outcomes. Nonetheless, the Constant-Curvature EKF ($M^{\mathit{(0)}}$) remains a widely accepted baseline[ataka2016real, loo2019non, peng2024tendon], and 1st-order polynomial curvature ($M^{\mathit{(1)}}$) aligns with popular linear-curvature assumptions (e.g., Euler curves[rao2022shape] or Clothoid curves[gandhi2023shape]). Hence, assessing our IMM-EKFs under identical conditions against these standard single-model EKFs demonstrates the measurable advantages and robustness offered by the proposed stochastic observer in polynomial curvature space. The ability to flexibly employ higher-order models ($M^{\mathit{(2)}}$) further underscores its potential for improved accuracy when the application demands richer shape representations.
@@ VIII Conclusion
An extended Supplementary Appendix (Appendices A–C) is available as an ancillary file with the arXiv version of this article[zhang2022stochastic].
@@ Acknowledgement
[Uncaptioned image] Guoqing Zhang received the B.E. degree in processing equipment and control engineering from Tianjin University, Tianjin, China, in 2014 and the M.E. degree in Chemical machinery from Xi’an Jiaotong University, Xi’an, China and the M.S. degree in mechanical engineering from Columbia University, New York, NY, USA, in 2019. He is currently a Ph.D. student studying surgical continuum robotics at Stevens Institute of Technology, Hoboken, NJ USA. His research interests include modeling, control, and estimation theory of continuum manipulators and other robotic systems.
[Uncaptioned image] Long Wang received the B.S. degree from Tsinghua University, Beijing, China, in 2010, the M.S. degree from Columbia University, New York, NY, USA, in 2012, and the Ph.D. degree from Vanderbilt University, Nashville, TN, USA, in 2019, all in mechanical engineering. He was a Postdoctoral Researcher with the Department of Mechanical Engineering, Columbia University, New York, NY, from 2018 to 2019. He is currently an Assistant Professor with the Department of Mechanical Engineering, Stevens Institute of Technology, Hoboken, NJ, USA. His research interests include modeling and calibration of continuum robots, surgical robotics, telemanipulation, manipulation, and grasping.